OpenStudy (anonymous):

A meteor streaking through the night sky is located with radar. At point its coordinates are (5.20 , 1.35), and 1.12 later its has moved to point with coordinates (6.24 , 0.925 ).

OpenStudy (anonymous):

Find the component of its average velocity between A and B. Find the component of its average velocity between A and B. Find the magnitude of its average velocity between these two points. Find the direction of its average velocity between these two points.

OpenStudy (jfraser):

you need units with those numbers, or nothing is going to help

OpenStudy (anonymous):

km/s for the first 3 @JFraser and the last one degrees counterwise from axis

OpenStudy (anonymous):

FIND THE DISPLACEMENT= R2 -R1

OpenStudy (anonymous):

= (6.24 , 0.925 ) -(5.20 , 1.35)

OpenStudy (anonymous):

=(1.04, -0.425)

OpenStudy (anonymous):

so av velocity=(1.04/1.12, -0.425/1.12)

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

yes thank you.

OpenStudy (anonymous):

then u can find mag of av velocity and direction of av velocity